3 quarts of water are needed to water 1 square foot of lawn. How many gallons of water will be need to water a lawn that measures 30 feets by 60 feet?
Found 2 solutions by richwmiller, acs1234:
Answer by richwmiller(17219)

the steps to solve this are:
1=find the area in square feet.
2- multiply by 3 to find gallons of water

Answer by acs1234(1)

First you would need to find how many square feet are in the lawn.
30*60=1800
Next, you know that there are four quarts in a gallon, and you need 3 quarts per square foot, so that means you need 3/4 of a gallon of water per square foot.
3 quarts
--------- = 3/4
4 quarts
Then, you would multiply the number of gallons you need per square foot by the number of square feet in the lawn.
3/4*1800=1350

1350 gallons of water will be needed to water a 30 by 60 ft lawn, watering 3 quarts of water per square foot.